OpenStudy (anonymous):

Find the derivative of the function. y = (5x − 3)^4(7x2 − 2)^−3 will fan and medal!!!

OpenStudy (anonymous):

product rule plus chain rule

OpenStudy (anonymous):

\[\left(fg\right)'=f'g+g'f\]with \[f(x)=(5x-3)^4, f'(x)=4(5x-3)^2\times 5\] \[g(x)=(7x^2-2)^{-3}, g'(x)=-3(7x^2-2)^{-4}\times 14x\]

OpenStudy (anonymous):

4(5x-3^2(5)(7x^2-2)^-3+-3(7x^2-2)^-4(14)(5x-3)^4 @satellite73